OpenStudy (anonymous):

what is x=5 if x is square rooted ?

jimthompson5910 (jim_thompson5910):

so the equation is \[\large \sqrt{x} = 5\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\sqrt{x} = x^{\frac{ 1 }{ 2 }}\] so square both sides to find x

OpenStudy (magbak):

\[\sqrt(x)=5\] \[(\sqrt(x)))^2=5^2\]

OpenStudy (magbak):

\[so x=25\]

OpenStudy (anonymous):

so...im terribly confused now haha.. id just just like the answer if you could

OpenStudy (magbak):

So x=25

OpenStudy (anonymous):

thank you!

OpenStudy (magbak):

You are welcome.